OpenStudy (anonymous):

What is factored version of b^2-16

jimthompson5910 (jim_thompson5910):

hint1: 16 is a perfect square; 16 = 4^2 hint2: a^2 - b^2 = (a-b)*(a+b) is the difference of squares rule

OpenStudy (anonymous):

B-4 ??

jimthompson5910 (jim_thompson5910):

that's part of it

OpenStudy (anonymous):

(B-4)(b-4) ??

jimthompson5910 (jim_thompson5910):

example: 9x^2 - 25 = (3x)^2 - 5^2 = (3x-5)(3x+5)

jimthompson5910 (jim_thompson5910):

(b-4)(b-4) is close but not quite there

OpenStudy (anonymous):

(B-4)(b+4)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

I recommend expanding that out to confirm the answer

OpenStudy (anonymous):

Thank you :) can you help with a couple more??

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Y^2-2y-15 Would it be (y-5)(y+3) ?

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

Okay :) what about 3x^2+9x?

jimthompson5910 (jim_thompson5910):

hint: factor out the GCF

OpenStudy (anonymous):

Would it be (x+3)^2

jimthompson5910 (jim_thompson5910):

hint: \[\Large 3x^2 + 9x\] \[\Large 3x*x + 3x*3\] \[\Large \color{red}{3x}*x + \color{red}{3x}*3\]

OpenStudy (anonymous):

X+3 ??

jimthompson5910 (jim_thompson5910):

do you see how there's the common factor of 3x? we can factor it out to get 3x(x+3) if you distribute the 3x back in, you should get the original expression back

OpenStudy (anonymous):

X^2+3

jimthompson5910 (jim_thompson5910):